# Einstein from Noise: Statistical Analysis

### v2

@@ 6 Discussion and outlook
In this work, we have shown that the Fourier phases of the EfN estimator converge to those of the template signal for an asymptotic number of observations. Since Fourier phases are crucial for perceiving image structure, the reconstructed image appears structurally similar to the template signal, even in cases where the estimator’s spectral magnitudes differ from those of the template [33, 41]. We have also shown that the Fourier phases of spectral components with higher magnitudes converge faster, leading to faster structural similarity in the overall image perception. In addition, we have extended our analysis beyond white Gaussian noise, examining other noise models. We have shown that the EfN estimator remains positively correlated with the template for arbitrary noise settings, and we have analyzed the Fourier phases convergence properties for high-dimensional i.i.d. noise (which is not necessarily Gaussian) and circulant Gaussian noise.
@@ 6.2 Future work
A natural direction for future work is to extend the EfN analysis beyond the simplified setting of cyclic translations, as defined in (3.1), to more general group actions, particularly those arising in practical applications such as cryo-EM. In this context, the relevant transformations are elements of the rotation group $\mathsf{SO}(3)$, and the postulated observations are two-dimensional projections of a three-dimensional structure rather than simple translations of a one-dimensional signal. However, extending the analysis to non-abelian groups presents more substantial challenges. In particular, the property of circular Gaussian statistics, which underpins the EfN analysis for cyclic groups, does not naturally extend to the non-abelian setting. Preliminary simulations for the non-abelian dihedral group (not shown here) indicate that the convergence of the EfN estimator’s Fourier phases observed in the abelian setting does not directly carry over. We suspect, however, that analogous spectral or structural biases may still arise, potentially governed by representation-theoretic properties of the group action.
@@ Acknowledgment
T.B. is supported in part by BSF under Grant 2020159, in part by NSF-BSF under Grant 2019752, and in part by ISF under Grant 1924/21. W.H. is supported by ISF under Grant 1734/21.
@@
Appendix A provides general preliminaries used throughout the paper, including notation and common technical tools. Appendix B presents the auxiliary lemmas required for Theorem 4.1, whose full proof appears in Appendix C. For Theorem 4.3, the necessary supporting results are given in Appendices D and E, with the proof provided in Appendix F. Appendix G contains the proof of Proposition 5.1, establishing the positive correlation property. Appendix H proves Theorem 5.2, which extends the results to high-dimensional settings with i.i.d. noise that is not necessarily Gaussian. Finally, Appendix I provides the proof of Proposition 5.4, addressing the case of structured noise with a circulant Gaussian covariance.
@@ Appendix A Preliminaries
Before we delve into the proofs of Theorems 4.1 and 4.3, we fix notations and definitions and prove auxiliary results that will be used in the proofs.
@@ A.1 Notations
for $10\leq r\leq d-$ 1, and therefore, $\mathsf{\hat{R}}_{i}=\operatorname*{arg\,max}_{0\leq r\leq{d-1}}\mathsf{S}_{i}[r]$. We note that for any $10\leq i\leq M-$ 1, the random vector $\mathsf{S}_{i}\triangleq(\mathsf{S}_{i}[0],\mathsf{S}_{i}[1],\ldots,\mathsf{S}_{i}[d-1])^{T}$ is Gaussian distributed, with zero mean vector, and a circulant covariance matrix; therefore, it is a cyclo-stationary random process.
Our goal is to investigate the phase and magnitude of the estimator $\hat{\mathsf{X}}$ in (3.7). Simple manipulations reveal that, for any $10\leq k\leq d-$ 1, the estimator’s phases are given by,
@@ A.2 The convergence of the Einstein from Noise estimator
${\mathsf{a.s.}}\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\cos\left(\phi_{e,1}[k]\right)\right]+j\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\sin\left(\phi_{e,1}[k]\right)\right],$ (A.13)
@@ A.3 Conditioning on the Fourier frequency noise component
$=\sigma^{2}\sum_{\ell=0}^{d-1}|\tilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}(r-s)\right),$ (A.20)
for $10\leq r,s\leq d-$ 1, where $\tilde{\mathsf{X}}_{k}$ is defined by:
$\tilde{\mathsf{X}}_{k}[\ell]\triangleq\begin{cases}0&\mathsf{if}\;\ell=k,d-k,\\ \mathsf{X}[\ell]&\mathsf{if}\;\ell=0,d/2,\\ \sqrt{2}\cdot\mathsf{X}[\ell]&\mathsf{otherwise}.\end{cases}$ (A.21)
@@
Note that the conditional process $\mathsf{S}_{i}|\mathsf{N}_{i}[k]$ is Gaussian because it is given by a linear transform of i.i.d. Gaussian variables. Also, since its covariance matrix is circulant and depends only on the difference between the two indices, i.e., $\mathsf{\Sigma}_{k,i}[r,s]=\sigma_{k,i}[\left|r-s\right|]$, it is cycle-stationary with a cosine trend. The eigenvalues of this circulant matrix are given by the DFT of its first row, and thus its $\ell$-th eigenvalue equals $|\tilde{\mathsf{X}}_{k}[\ell]|^{2}$, for $10\leq\ell\leq d-$ 1.
@@
where we have used the property of $\mathsf{X}[k]=\overline{\mathsf{X}[d-k]},\ \mathsf{N}_{i}[k]=\overline{\mathsf{N}_{i}[d-k]}$, (A.1). Clearly, as $\mathbb{E}\left[\mathsf{N}_{i}\left[\ell\right]\right]=0$, for every $10\leq\ell\leq d-$ 1, we have,
@@
for every $10\leq\ell\leq d-$ 1. Combining (A.22) and (A.23) results,
@@
=$\mathcal{I}=\left\{1,2,\dots k-1,k+1,\dots,d/2-1\right\},$ (A.27)
@@
As the sequences $\left\{|\mathsf{N}_{i}[\ell]|\right\}_{\ell=0}^{d/2}$ and $\left\{\phi_{\mathsf{N}_{i}}[\ell]\right\}_{\ell=0}^{d/2}$ are two independent sequences of i.i.d. random variables, that satisfy $\mathsf{N}_{i}[\ell]=\overline{\mathsf{N}_{i}[d-\ell]}$, as well as $\mathsf{X}[\ell]=\overline{\mathsf{X}[d-\ell]}$, we have,
@@
Since the terms in the sum on the right-hand side of (A.28) are independent, that is, $\mathbb{E}\left[\mathsf{N}_{i}\left[\ell_{1}\right]\mathsf{N}_{i}\left[\ell_{2}\right]\right]=\mathbb{E}\left[\left|\mathsf{N}_{i}\left[\ell_{1}\right]\right|^{2}\right]\delta_{\ell_{1},\ell_{2}}$, it follows that,
@@
$=\cos\left(\frac{2\pi\ell(r-s)}{d}\right)+\cos\left(\frac{2\pi\ell(r+s)}{d}+2\left(\phi_{\mathsf{N}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)\right).$ (A.30)
For every $10\leq r,s\leq d-$ 1, since the sequences $\left\{|\mathsf{N}_{i}[\ell]|\right\}_{\ell=0}^{d/2}$ and $\left\{\phi_{\mathsf{N}_{i}}[\ell]\right\}_{\ell=0}^{d/2}$ are independent and consist of i.i.d. random variables, with $\mathbb{E}\left[\left|\mathsf{N}_{i}[k]\right|^{2}\right]=\sigma^{2}$ and phases $\phi_{\mathsf{N}_{i}}[k]$ uniformly distributed over $[-\pi,\pi)$, and by applying the trigonometric identity (A.30), it follows that,
@@
$=\mathbb{E}\left[\left|\mathsf{N}_{i}[\ell]\right|^{2}\right]\cos\left(\frac{2\pi\ell(r-s)}{d}\right)=\sigma^{2}\cos\left(\frac{2\pi\ell(r-s)}{d}\right).$ (A.31)
@@
$\frac{1}{\sigma^{2}}\mathbb{E}$ $\left[\rho_{k,i}\left[r\right]\rho_{k,i}\left[s\right]|\mathsf{N}_{i}[k]\right]={{\sum_{\ell\in\left\{0,d/2\right\}}{\left|\mathsf{X}[\ell]\right|^{2}\,\cos\left(\frac{2\pi\ell}{d}(r-s)\right)}}}$
@@
$\mathbb{E}$ $\left[\rho_{k,i}\left[r\right]\rho_{k,i}\left[s\right]|\mathsf{N}_{i}[k]\right]=\sigma^{2}\sum_{\ell=0}^{d-1}|\tilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}(r-s)\right),$
for $\tilde{\mathsf{X}}_{k}[\ell]$ defined in (A.21), which complete the proof. ∎
A.4 Auxiliary to Proposition B.2

### v3

@@ 6 Discussion and outlook
In this work, we have shown that the Fourier phases of the EfN estimator converge to those of the template signal for an asymptotic number of observations. Since Fourier phases are crucial for perceiving image structure, the reconstructed image appears structurally similar to the template signal, even in cases where the estimator’s spectral magnitudes differ from those of the template [36, 45]. We have also shown that the Fourier phases of spectral components with stronger magnitudes converge faster, leading to faster structural similarity in the overall image perception. In addition, we have extended our analysis beyond white Gaussian noise, examining other noise models. We have shown that the EfN estimator remains positively correlated with the template for arbitrary noise settings, and we have analyzed the Fourier phases convergence properties for high-dimensional i.i.d. noise (which is not necessarily Gaussian) and circulant Gaussian noise.
@@ 6.2 Future work
A natural direction for future work is to extend the EfN analysis beyond the simplified setting of cyclic translations, as defined in (2.1), to more general group actions, particularly those arising in practical applications such as cryo-EM. In this context, the relevant transformations are elements of the rotation group $\mathsf{SO}(3)$, and the postulated observations are two-dimensional projections of a three-dimensional structure rather than simple translations of a one-dimensional signal. Recent empirical evidence suggests that template-induced bias can persist in such non-abelian settings as well, including under $\mathsf{SO}(3)$ group actions [57].
However, extending the theoretical analysis to non-abelian groups presents more substantial challenges. In particular, the property of circular Gaussian statistics, which underpins the EfN analysis for cyclic groups, does not naturally extend to the non-abelian setting. Preliminary simulations for the non-abelian dihedral group (not shown here) indicate that the convergence of the EfN estimator’s Fourier phases observed in the abelian setting does not directly carry over. We nonetheless expect that alignment over a richer group may still induce systematic structural biases, potentially governed by the group’s representation-theoretic decomposition. This broader perspective is consistent with recent observations of confirmation bias effects even in unstructured latent-variable models: for Gaussian mixture models with pure-noise data, a single iteration of population $k$-means or EM initialized at hypothesized centroids produces estimates that remain positively correlated with the initialization [5], suggesting that algorithmic bias can persist beyond group-alignment settings, albeit often in weaker forms than the phase-locking phenomenon characterized here.
@@ Acknowledgment
T.B. is supported in part by BSF under Grant 2020159, in part by NSF-BSF under Grant 2024791, and in part by ISF under Grant 1924/21. W.H. is supported by ISF under Grant 1734/21.
@@
Appendix A provides general preliminaries used throughout the paper, including notation and common technical tools. Appendix B presents the auxiliary lemmas required for Theorem 4.1, whose full proof appears in Appendix B.5. For Theorem 4.3, the necessary supporting results are given in Appendices C, with the proof provided in Appendix D. Appendix E contains the proof of Proposition 5.1, establishing the positive correlation property. Appendix F proves Theorem 5.2, which extends the results to high-dimensional settings with i.i.d. noise that is not necessarily Gaussian. Finally, Appendix G provides the proof of Proposition 5.4, addressing the case of structured noise with a circulant Gaussian covariance.
@@ Appendix A Preliminaries
Before we delve into the proofs, we fix notations and definitions and prove auxiliary results that will be used in the proofs.
@@ A.1 Notations
for $0\leq r\leq d-1$, and therefore, $\mathsf{\hat{R}}_{i}=\operatorname*{arg\,max}_{0\leq r\leq{d-1}}\mathsf{S}_{i}[r]$. We note that for any $0\leq i\leq M-1$, the random vector $\mathsf{S}_{i}\triangleq(\mathsf{S}_{i}[0],\mathsf{S}_{i}[1],\ldots,\mathsf{S}_{i}[d-1])^{T}$ is Gaussian distributed, with zero mean vector, and a circulant covariance matrix; therefore, it is a cyclo-stationary random process.
Our goal is to investigate the phase and magnitude of the estimator $\hat{\mathsf{X}}$ in (2.7). Simple manipulations reveal that, for any $0\leq k\leq d-1$, the estimator’s phases are given by,
@@ A.2 The convergence of the Einstein from Noise estimator
$\xrightarrow[]{\mathsf{a.s.}}\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\cos\left(\phi_{e,1}[k]\right)\right]+j\mathbb{E}\left[\left|\mathsf{N}_{1}[k]\right|\sin\left(\phi_{e,1}[k]\right)\right],$ (A.13)
@@ A.3 Conditioning on the Fourier frequency noise component
$=\frac{\sigma^{2}}{2}\sum_{\ell=0}^{d-1}|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}(r-s)\right),$ (A.20)
for $0\leq r,s\leq d-1$, where $\widetilde{\mathsf{X}}_{k}$ is defined by:
$\widetilde{\mathsf{X}}_{k}[\ell]\triangleq\begin{cases}0&\mathsf{if}\ell=k,d-k,\\ \mathsf{X}[\ell]&\mathsf{if}\ell=0,d/2,\\ \sqrt{2}\cdot\mathsf{X}[\ell]&\mathsf{otherwise}.\end{cases}$ (A.21)
@@
Note that the conditional process $\mathsf{S}_{i}|\mathsf{N}_{i}[k]$ is Gaussian because it is given by a linear transform of i.i.d. Gaussian variables. Also, since its covariance matrix is circulant and depends only on the difference between the two indices, i.e., $\mathsf{\Sigma}_{k,i}[r,s]=\sigma_{k,i}[\left|r-s\right|]$, it is cycle-stationary with a cosine trend. The eigenvalues of this circulant matrix are given by the DFT of its first row, and thus its $\ell$-th eigenvalue equals $|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}$, for $0\leq\ell\leq d-1$.
@@
where we have used the property of $\mathsf{X}[k]=\overline{\mathsf{X}[d-k]},\ \mathsf{N}_{i}[k]=\overline{\mathsf{N}_{i}[d-k]}$, (A.1). Clearly, as $\mathbb{E}\left[\mathsf{N}_{i}\left[\ell\right]\right]=0$, for every $0\leq\ell\leq d-1$, we have,
@@
for every $0\leq\ell\leq d-1$. Combining (A.22) and (A.23) results,
@@
$\mathcal{I}=\left\{1,2,\dots k-1,k+1,\dots,d/2-1\right\},$ (A.27)
@@
As the sequences $\left\{|\mathsf{N}_{i}[\ell]|\right\}_{\ell=0}^{d/2}$ and $\left\{\phi_{\mathsf{N}_{i}}[\ell]\right\}_{\ell=0}^{d/2}$ satisfy $\mathsf{N}_{i}[\ell]=\overline{\mathsf{N}_{i}[d-\ell]}$, as well as $\mathsf{X}[\ell]=\overline{\mathsf{X}[d-\ell]}$, we have,
@@
Since the terms in the sum on the right-hand side of (A.28) are independent, that is, $\mathbb{E}\left[\mathsf{N}_{i}\left[\ell_{1}\right]\overline{\mathsf{N}_{i}\left[\ell_{2}\right]}\right]=\mathbb{E}\left[\left|\mathsf{N}_{i}\left[\ell_{1}\right]\right|^{2}\right]\delta_{\ell_{1},\ell_{2}}$, it follows that,
@@
$=\frac{1}{2}\cos\left(\frac{2\pi\ell(r-s)}{d}\right)+\frac{1}{2}\cos\left(\frac{2\pi\ell(r+s)}{d}+2\left(\phi_{\mathsf{N}_{i}}[\ell]-\phi_{\mathsf{X}}[\ell]\right)\right).$ (A.30)
For $\ell\in\left\{1,\ldots,\frac{d}{2}-1\right\}$, the DFT coefficients $N_{i}[\ell]$ are i.i.d. circular complex Gaussian (under white Gaussian noise in the time domain), hence independent across $\ell$, with $\phi_{N_{i}}[\ell]\sim\mathrm{Unif}\!\left[-\pi,\pi\right)$ independent of $\lvert N_{i}[\ell]\rvert$. Thus,
@@
$=\frac{1}{2}\mathbb{E}\left[\left|\mathsf{N}_{i}[\ell]\right|^{2}\right]\cos\left(\frac{2\pi\ell(r-s)}{d}\right)=\frac{\sigma^{2}}{2}\cos\left(\frac{2\pi\ell(r-s)}{d}\right).$ (A.31)
@@
$\frac{2}{\sigma^{2}}\mathbb{E}$ $\left[\rho_{k,i}\left[r\right]\rho_{k,i}\left[s\right]|\mathsf{N}_{i}[k]\right]={{\sum_{\ell\in\left\{0,d/2\right\}}{\left|\mathsf{X}[\ell]\right|^{2}\,\cos\left(\frac{2\pi\ell}{d}(r-s)\right)}}}$
@@
$\mathbb{E}$ $\left[\rho_{k,i}\left[r\right]\rho_{k,i}\left[s\right]|\mathsf{N}_{i}[k]\right]=\frac{\sigma^{2}}{2}\sum_{\ell=0}^{d-1}|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}(r-s)\right),$
for $\widetilde{\mathsf{X}}_{k}[\ell]$ defined in (A.21), which completes the proof. ∎
A.4 Uniqueness of the maximizer
Lemma A.3 (Uniqueness of the maximizer).
Recall the definition of $\mathsf{S}_{i}$ from (A.5). Assume $d\geq 6$ is even. Fix $k\in\{1,\ldots,\frac{d}{2}-1,\frac{d}{2}+1,\ldots,d-1\}$ and recall from Lemma A.1 that
@@ A.4 Uniqueness of the maximizer
$\mathsf{S}_{i}|\mathsf{N}_{i}[k]\sim\mathcal{N}(\mu_{k,i},\Sigma_{k,i}),\qquad\Sigma_{k,i}[r,s]=\frac{\sigma^{2}}{2}\sum_{\ell=0}^{d-1}\big|\widetilde{\mathsf{X}}_{k}[\ell]\big|^{2}\cos\!\left(\frac{2\pi\ell}{d}(r-s)\right).$ (A.34)
Assume moreover that the template spectrum is non-vanishing, i.e. $\mathsf{X}[\ell]\neq 0$ for all $\ell\in\{0,\ldots,d-1\}$. Then, for every $r\neq s$,
$\operatorname{Var}\!\Big((\mathsf{S}_{i}|\mathsf{N}_{i}[k])_{r}-(\mathsf{S}_{i}|\mathsf{N}_{i}[k])_{s}\Big)$ $=\sigma^{2}\sum_{\ell=0}^{d-1}\big|\widetilde{\mathsf{X}}_{k}[\ell]\big|^{2}\Big(1-\cos\!\big(\tfrac{2\pi\ell}{d}(r-s)\big)\Big)>0.$ (A.35)
Consequently, the maximizer $\mathsf{\hat{R}}_{i}=\operatorname*{arg\,max}_{0\leq r\leq d-1}(\mathsf{S}_{i}|\mathsf{N}_{i}[k])_{r}$ is unique almost surely.
Proof of Lemma A.3.
Fix $r\neq s$ and set $m\triangleq r-s\not\equiv 0\ (\mathrm{mod}\ d)$. Using the covariance formula and circulantness,
$\operatorname{Var}\!\Big((\mathsf{S}_{i}|\mathsf{N}_{i}[k])_{r}-(\mathsf{S}_{i}|\mathsf{N}_{i}[k])_{s}\Big)$ $=\Sigma_{k,i}[r,r]+\Sigma_{k,i}[s,s]-2\Sigma_{k,i}[r,s]$
$=\sigma^{2}\sum_{\ell=0}^{d-1}\big|\widetilde{\mathsf{X}}_{k}[\ell]\big|^{2}\Big(1-\cos\!\big(\tfrac{2\pi\ell}{d}m\big)\Big),$ (A.36)
which proves the identity in (A.35). Each summand is nonnegative.
It remains to show strict positivity. We show that there is at least one term in the sum that is strictly positive. Define
$H_{m}\triangleq\{\ell\in\{0,\ldots,d-1\}:\ \ell m\equiv 0\ (\mathrm{mod}\ d)\}.$ (A.37)
Then $|H_{m}|=\gcd(d,m)\leq d/2$ since $m\not\equiv 0\ (\mathrm{mod}\ d)$. Because $d\geq 6$, we have
$\big|\{0,\ldots,d-1\}\setminus(H_{m}\cup\{k,d-k\})\big|$ $\geq d-|H_{m}|-2$ (A.38)
$\geq d-\frac{d}{2}-2=\frac{d}{2}-2>0.$ (A.39)
Hence we may choose $\ell_{0}\in\{0,\ldots,d-1\}\setminus(H_{m}\cup\{k,d-k\})$. By construction, $\ell_{0}\notin H_{m}$ implies $\ell_{0}m\not\equiv 0\ (\mathrm{mod}\ d)$, i.e. $\cos(2\pi\ell_{0}m/d)\neq 1$ and thus $1-\cos(2\pi\ell_{0}m/d)>0$. Moreover, since $\widetilde{\mathsf{X}}_{k}[\ell]=0$ only for $\ell\in\{k,d-k\}$ and $\mathsf{X}[\ell]\neq 0$ for all $\ell$, we have $|\widetilde{\mathsf{X}}_{k}[\ell_{0}]|^{2}>0$. Therefore the $\ell_{0}$-term in the sum is strictly positive, and the entire variance is strictly positive, proving (A.35).
Finally, for any $r\neq s$, the difference $(\mathsf{S}_{i}|\mathsf{N}_{i}[k])_{r}-(\mathsf{S}_{i}|\mathsf{N}_{i}[k])_{s}$ is a non-degenerate Gaussian, hence $\mathbb{P}\!\big((\mathsf{S}_{i}|\mathsf{N}_{i}[k])_{r}=(\mathsf{S}_{i}|\mathsf{N}_{i}[k])_{s}\big)=0$. A union bound over finitely many pairs implies ties occur with probability 0, so the maximizer is unique almost surely. ∎
A.5 Positive probability of each maximizer event
Lemma A.4 (Positive probability of each strict maximizer event).
Recall the definition of $\mathsf{S}_{i}$ from (A.5). Assume $d\geq 6$ is even. Fix $k\in\{1,\ldots,\frac{d}{2}-1,\frac{d}{2}+1,\ldots,d-1\}$ and recall from Lemma A.1 that
$\mathsf{S}_{i}|\mathsf{N}_{i}[k]\sim\mathcal{N}(\mu_{k,i},\Sigma_{k,i}),\qquad\Sigma_{k,i}[r,s]=\frac{\sigma^{2}}{2}\sum_{\ell=0}^{d-1}\big|\widetilde{\mathsf{X}}_{k}[\ell]\big|^{2}\cos\!\Big(\frac{2\pi\ell}{d}(r-s)\Big).$ (A.40)
Assume moreover that the template spectrum is non-vanishing, i.e. $\mathsf{X}[\ell]\neq 0$ for all $\ell\in\{0,\ldots,d-1\}$. Then, for every $r\in\{0,\ldots,d-1\}$, the event
$\mathcal{C}_{r}\triangleq\Big\{\,(\mathsf{S}_{i}|\mathsf{N}_{i}[k])_{r}>\max_{t\neq r}(\mathsf{S}_{i}|\mathsf{N}_{i}[k])_{t}\,\Big\}$ (A.41)
has strictly positive probability:
$\mathbb{P}\big(\mathcal{C}_{r}\,\big|\,\mathsf{N}_{i}[k]\big)>0,$ (A.42)
for almost every realization of $\mathsf{N}_{i}[k]$.
Proof of Lemma A.4.
Fix $k$ as in the statement and condition on a realization of $\mathsf{N}_{i}[k]$. Then $Y\triangleq\mathsf{S}_{i}|\mathsf{N}_{i}[k]\sim\mathcal{N}(m,\Sigma)$ with $m=\mu_{k,i}$ and $\Sigma=\Sigma_{k,i}$. By Lemma A.3, for every $r\neq s$, $\operatorname{Var}(Y_{r}-Y_{s})>0$, hence $Y_{r}-Y_{s}$ is a non-degenerate Gaussian and $\mathbb{P}(Y_{r}=Y_{s})=0$. In particular, ties occur with probability 0.
Since $\Sigma$ is real, symmetric, and circulant, it is diagonalized by the DFT basis: there exist eigenvectors $\{f_{\ell}\}_{\ell=0}^{d-1}$ (the Fourier modes) such that the corresponding eigenvalues $\{\lambda_{\ell}\}_{\ell=0}^{d-1}$ are given by the DFT of the first row of $\Sigma$. In Model A.40, Lemma A.1 shows that these eigenvalues satisfy
$\lambda_{\ell}\ \propto\ |\widetilde{\mathsf{X}}_{k}[\ell]|^{2},\qquad\ell=0,\ldots,d-1.$ (A.43)
Under the non-vanishing spectrum assumption $\mathsf{X}[\ell]\neq 0$ for all $\ell$, and by the definition (A.21), we have
$|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}>0\ \quad\text{for all }\ell\notin\{k,d-k\},$ (A.44)
while $|\widetilde{\mathsf{X}}_{k}[k]|^{2}=|\widetilde{\mathsf{X}}_{k}[d-k]|^{2}=0$. Hence $\Sigma$ has exactly two zero eigenvalues, corresponding to the $\ell=k$ and $\ell=d-k$ Fourier modes. In the real domain, these two modes are equivalently represented by the cosine and sine vectors $c,s\in\mathbb{R}^{d}$
$c_{t}=\cos\!\Big(\tfrac{2\pi k}{d}t\Big),\qquad s_{t}=\sin\!\Big(\tfrac{2\pi k}{d}t\Big),$ (A.45)
for $t=0,\ldots,d-1$, so
$\ker(\Sigma)=\operatorname{span}\{c,s\}.$ (A.46)
Let $L\triangleq\operatorname{range}(\Sigma)$. Since $\Sigma$ is symmetric, we have $L=(\ker\Sigma)^{\perp}$, and therefore the Gaussian vector $Y\sim\mathcal{N}(m,\Sigma)$ is supported on the affine subspace $m+L$.
We fix $r\in\{0,\ldots,d-1\}$ and consider the open cone
$\mathcal{C}_{r}\triangleq\{y\in\mathbb{R}^{d}:\ y_{r}>\max_{t\neq r}y_{t}\}.$ (A.47)
We claim that $\mathcal{C}_{r}\cap(m+L)\neq\emptyset$, i.e., the affine support contains at least one point whose unique largest coordinate is the $r$-th.
To build such a point, we start from the $r$-th standard basis vector $u=e_{r}$, whose maximizer is trivially at $r$ (i.e., $e_{r}\in\mathcal{C}_{r}$), but which may fail to belong to $m+L$. We therefore remove from $u$ its components along the two forbidden directions $c$ and $s$ (which span $\ker(\Sigma)$ (A.46)). We define
$\alpha\triangleq\frac{\langle u,c\rangle}{\langle c,c\rangle},\qquad\beta\triangleq\frac{\langle u,s\rangle}{\langle s,s\rangle},\qquad v\triangleq u-\alpha\,c-\beta\,s.$ (A.48)
Because $k\notin\{0,d/2\}$, the sine and cosine vectors are orthogonal and have equal energy:
$\langle c,s\rangle=0,\qquad\langle c,c\rangle=\langle s,s\rangle=d/2.$ (A.49)
Hence $\langle v,c\rangle=\langle v,s\rangle=0$, which means $v\in(\ker\Sigma)^{\perp}=L$.
Next we show that $v$ remains strongly peaked at the $r$-th coordinate. Since $u=e_{r}$, we have $\langle u,c\rangle=c_{r}$ and $\langle u,s\rangle=s_{r}$, and therefore
$v_{r}=1-\frac{c_{r}^{2}}{\langle c,c\rangle}-\frac{s_{r}^{2}}{\langle s,s\rangle}=1-\frac{c_{r}^{2}+s_{r}^{2}}{d/2}=1-\frac{2}{d},$ (A.50)
where we used $c_{r}^{2}+s_{r}^{2}=1$ (A.45). For $t\neq r$, since $u=e_{r}$ is the $r$-th standard basis vector, we have $u_{t}=0$; thus, using (A.45) and (A.49), we obtain
$v_{t}=-\alpha\,c_{t}-\beta\,s_{t}=-\frac{c_{r}c_{t}+s_{r}s_{t}}{d/2}=-\frac{2}{d}\cos\!\Big(\tfrac{2\pi k}{d}(t-r)\Big),$ (A.51)
so $|v_{t}|\leq 2/d$ for all $t\neq r$. Consequently, letting
$\delta\triangleq v_{r}-\max_{t\neq r}v_{t},$ (A.52)
we obtain the uniform lower bound for $d>4$
$\delta\geq\Big(1-\frac{2}{d}\Big)-\frac{2}{d}=1-\frac{4}{d}>0.$ (A.53)
Thus, within the direction $v\in L$, the $r$-th coordinate exceeds every other coordinate by at least $\delta$.
Finally, because the affine support is $m+L$, any point of the form $y(\gamma)=m+\gamma v$ lies in the support. The offset $m$ may change the ordering for small $\gamma$, but scaling $\gamma$ makes the $v$-term dominate. In particular, let
$M\triangleq\max_{t\neq r}|m_{r}-m_{t}|.$ (A.54)
Then for any $t\neq r$,
$y_{r}(\gamma)-y_{t}(\gamma)=(m_{r}-m_{t})+\gamma(v_{r}-v_{t})\ \geq\ -M+\gamma\delta.$ (A.55)
Choosing $\gamma>M/\delta$ guarantees $y_{r}(\gamma)>y_{t}(\gamma)$ for all $t\neq r$, hence
$y(\gamma)\in\mathcal{C}_{r}\cap(m+L),$ (A.56)
proving $\mathcal{C}_{r}\cap(m+L)\neq\emptyset$.
Since $\mathcal{C}_{r}$ is open in $\mathbb{R}^{d}$, the intersection $\mathcal{C}_{r}\cap(m+L)$ contains a nonempty open subset of the affine support $m+L$. A Gaussian measure assigns positive probability to any nonempty open subset of its affine support [26]; therefore, $\mathbb{P}(Y\in\mathcal{C}_{r})>0$, or equivalently $\mathbb{P}\!\big(\mathcal{C}_{r}\,\big|\,\mathsf{N}_{i}[k]\big)>0$, completing the proof. ∎
A.6 Auxiliary result for Proposition B.2